OpenStudy (anonymous):

a wheel makes 3 13/20 revolutions per minute. if it rotates for 44 minutes how many revolutions does it make?

OpenStudy (p0sitr0n):

3 13/20 = 73/20 73/20 * 44 =160.6 so 160 revolutions

OpenStudy (anonymous):

thank you so much